OpenStudy (anonymous):

What is the derivative of (-x^2)(e^(x/2))+1 Step by step please :)

OpenStudy (anonymous):

\[-x^2(e^{\frac{1}{2}}+1)\] is either a trick question or a typo

OpenStudy (anonymous):

\(e^{\frac{1}{2}}+1\) is a constant. is there an \(x\) involved some where in that part?

OpenStudy (anonymous):

\[(-x ^{2} * e ^{1/2}) + 1\]

OpenStudy (anonymous):

You have the parentheses in the wrong place satellite, the '+1' does not go in there with the e

OpenStudy (anonymous):

but it is \(e^{\frac{1}{2}}=\sqrt{e}\) right? not \(e^{\frac{1}{2}x}\) for example

OpenStudy (anonymous):

the derivate of 1 is zero the derivate de polinomio for constante is the derivate of polinomio for te constant constant=\[-e ^{1/2}\]

OpenStudy (anonymous):

then the derivative is \(-2\sqrt{e}x\) and that is all

OpenStudy (anonymous):

Right, but the one is just there as a distraction, per se

OpenStudy (anonymous):

since \(e^{\frac{1}{2}}\) is just a number

OpenStudy (anonymous):

then derivate is \[2*x*(-e ^{1/2}) + 0\]

OpenStudy (anonymous):

or \[-2xe ^{1/2}\]

OpenStudy (anonymous):

Ankolix are you sure??

OpenStudy (anonymous):

yeah! only if the equation that you put was \[x ^{2}*(e ^{1/2})*(-1) + 1\]

OpenStudy (anonymous):

Yes that's right. Do you know how to get the critical values? I have x=o as one, but I need the second one

OpenStudy (anonymous):

when the derivate is 0... so, i think that only answer is x=0

OpenStudy (anonymous):

Well... My Teacher said that I am missing a critical value... I'm so confused :(

OpenStudy (anonymous):

can you write the exact original problem? the only place where \(-2\sqrt{e}x=0\) is if \(x=0\) i get the feeling something is missing

OpenStudy (anonymous):

perhaps it was something like \[-x^2e^{\frac{x}{2}}\] not to second guess you

OpenStudy (anonymous):

yeah, im agree with satellite73

OpenStudy (anonymous):

I promise you that the equation is (-x^2)(e^(1/2)) +1

OpenStudy (anonymous):

It's completely possible that he made a mistake when grading my paper

OpenStudy (anonymous):

Satellite my mistake, it is e^x/2

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

got logged out

OpenStudy (anonymous):

@satellite73 sorry to bother anyone here, if you have a chance could you possibly help me with my problem when you are done here? thanks.

OpenStudy (anonymous):

\[f(x)=-x^2e^{\frac{x}{2}}+1\] you need the product rule for the derivative

OpenStudy (anonymous):

\[f'(x)=-2xe^{\frac{x}{2}}-\frac{1}{2}x^2e^{\frac{x}{2}}\]

OpenStudy (anonymous):

factor out the \(xe^{\frac{x}{2}}\) and then set equal to zero

OpenStudy (anonymous):

\[f'(x)=-xe^{\frac{x}{2}}\left(2+\frac{1}{2}x\right)\]

OpenStudy (anonymous):

i see, i see..... ok Thank you this helped a lot :)

OpenStudy (anonymous):

one zero is if \(x=0\) the other if \(2+\frac{1}{2}x=0\) etc

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@erin i don't see your problem

OpenStudy (anonymous):

@satellite73 i'll repost it now